OpenStudy (n0o0sha):

Hi guys, please help me s=2πr^2+2πrh solve for h

OpenStudy (n0o0sha):

@YanaSidlinskiy

OpenStudy (will.h):

\[S = 2\pi r^2 + 2\pi rh\] is that the question?

OpenStudy (n0o0sha):

yup @Will.H

OpenStudy (n0o0sha):

can you help? @Will.H

OpenStudy (will.h):

yes am just trying to figure a way to remove r^2

OpenStudy (n0o0sha):

oh okay take your time lol @Will.H thank you in advance i really appreciate it :))

OpenStudy (will.h):

i made a mistake 1 moment lol i am too sleepy So i don't know where i got that 2 from here's the correct solve \[S = 2\pi r^2 + 2\pi rh\] \[S - 2\pi r^2 = 2\pi rh\] \[\frac{ S - 2\pi r^2 }{ 2\pi r } = h\]

OpenStudy (will.h):

Sorry for the confusion

OpenStudy (n0o0sha):

Its okay don't worry about it :) @Will.H

OpenStudy (will.h):

There's another solution if you ever heard about factoring here it is \[S = 2\pi r^2 + 2\pi rh\] There's a GCF in the right side of 2pi r \[S = 2\pi r(r + h)\] Now divide both sides by 2pi r \[\frac{ S }{ 2\pi r } = r + h\] and now we may subtract r from both sides to isolate h \[\frac{ S }{ 2\pi r } - r = h\] Personally i would prefer this way

jhonyy9 (jhonyy9):

@Will.H this is very very nice great job really

OpenStudy (will.h):

Thank you Jhony :) @jhonyy9

jhonyy9 (jhonyy9):

np welcome

OpenStudy (skullpatrol):

You could also use $$C=2\pi r$$ to make it look even simpler.